# The general position number under vertex and edge removal

### v1

@@ 6 Edge removing in general graphs
In this section we consider how much the general position number can be affected by removing an edge. Contrary to the vertex removal, we can give general sharp lower and upper bounds. To prove them, we first recall the following well-known sets from metric graph theory. For instance, these sets are constitutional stones of partial cubes [22], of distance-balanced graphs [10], of $\ell$-distance-balanced graphs [20], and of the Mostar index [1].
@@
Let $e=uv$ be an edge in a graph $G$ and let $x,y\in W_{uv}\cup\,_{v}W_{u}$. Then ${\cal P}_{G}(x,y)={\cal P}_{G-e}(x,y)$.
@@
Proof. Let $e=uv$ and let $X$ be an arbitrary ${\rm gp}$-set of $G$. Assume first that $u,v\notin X$. Then we claim that $X$ is also a general position set in $G-e$. To see it, consider any two vertices $x,y\in X$. Since ${\rm diam}(G)=2$ and $u,v\notin X$, no shortest $x,y$-path of $G$ contains the edge $uv$, hence ${\cal P}_{G}(x,y)={\cal P}_{G-e}(x,y)$. Thus $X$ is a general position set of $G-e$, so that in this case ${\rm gp}(G-e)\geq{\rm gp}(G)$. Assume second that $u\in X$. Then we claim that $X^{\prime}=X\setminus\{u\}$ is a general position set in $G-e$. If $v\notin X$, then we see by the same argument as above that $X^{\prime}$ is a general position set of $G-e$. It remains to consider the subcase when $v\in X$. Since ${\rm diam}(G)=2$, the only way that $X^{\prime}$ is not a general position set in $G-e$ would be when there is a shortest path $Q$ in $G-e$ containing $v$ and two other vertices of $X^{\prime}$. Since $Q$ is not a shortest path in $G$, there is a shortest path in $G$ containing $u$, $v$, and another vertex of $X$, a contradiction. We can conclude that $X^{\prime}$ is a general position set of $G-e$ and therefore ${\rm gp}(G-e)\geq{\rm gp}(G)-1$.
Let $Y$ be an arbitrary ${\rm gp}$-set of $G-e$. Assume that $u,v\notin Y$. We claim that $Y$ is a general position set of $G$. Consider any two vertices $x,y$ from $Y$ and let $P$ be an arbitrary shortest $x,y$-path in $G$. Then the path $P$ does not contain the edge $uv$ in $G$ as ${\rm diam}(G)=2$, hence ${\cal P}_{G}(x,y)={\cal P}_{G-e}(x,y)$. We thus have that $Y$ is a general position set of $G$, and ${\rm gp}(G)\geq{\rm gp}(G-e)$. Assume now that $u\in Y$. Then we claim that $Y^{\prime}=Y\setminus\{u\}$ is a general position set of $G-e$. If $v\notin Y$, then similarly as above, $Y^{\prime}$ is a general position set of $G$. It remains to consider the case when $v\in Y$. Suppose on the contrary that there are three vertices from $Y^{\prime}$ such that they lie on a common shortest path $P^{\prime}$ in $G$. Then the path $P^{\prime}$ contains $v$ and two other vertices from $Y^{\prime}$. Since $P^{\prime}$ is not a shortest path in $G-e$ and ${\rm diam}(G)=2$, there is a shortest path in $G$ containing $u$, $v$, and two another vertices of $Y^{\prime}$. It contradicts with the fact that ${\rm diam}(G)=2$. Hence $Y^{\prime}$ is a general position set of $G$ and we can conclude that ${\rm gp}(G)\geq{\rm gp}(G-e)-1$.
To show that the lower bound is sharp, consider once fan graphs $F_{n}$, this time of order $n=3k$, $k\geq 3$, see Fig. 6.
@@
Figure 6: Fan graph $F_{n}$ with $n=3k$.
By Proposition 2.4 we known that ${\rm gp}(F_{n})=\lceil\frac{2n}{3}\rceil$, see Fig. 6 where the black vertices form a ${\rm gp}$-set of $F_{n}$. Let $uv$ be the edge of $F_{n}$ as shown in the figure. Then it is straightforward to check that all the black vertices but $v$ form a ${\rm gp}$-set of $F_{n}-e$, hence ${\rm gp}(F_{n}-e)={\rm gp}(F_{n})-1$.
@@ Acknowledgements
This work was supported by the Slovenian Research Agency ARIS (research core funding P1-0297 and projects N1-0285, N1-0218).

### v2

@@ 6 Edge removing in general graphs
In this section we consider how much the general position number can be affected by removing an edge. Contrary to the vertex removal, we can give general sharp lower and upper bounds. To prove them, we first recall the following well-known sets from metric graph theory. For instance, these sets are of interest in the study of partial cubes [22], of distance-balanced graphs [10], of $\ell$-distance-balanced graphs [20], and of the Mostar index [1].
@@
Let $e=uv$ be an edge in a graph $G$ and let $x,y\in W_{uv}\cup\,_{v}W_{u}$. Then ${\cal P}_{G}(x,y)={\cal P}_{G-e}(x,y)$. In particular, $d_{G}(x,y)=d_{G-e}(x,y)$.
@@
Proof. Let $e=uv$ and let $X$ be an arbitrary ${\rm gp}$-set of $G$. Assume first that $u,v\notin X$. Then we claim that $X$ is also a general position set in $G-e$. To see it, consider any two vertices $x,y\in X$. Since ${\rm diam}(G)=2$ and $u,v\notin X$, no shortest $x,y$-path of $G$ contains the edge $uv$, hence ${\cal P}_{G}(x,y)={\cal P}_{G-e}(x,y)$. Thus $X$ is a general position set of $G-e$, so that in this case ${\rm gp}(G-e)\geq{\rm gp}(G)$. Assume second that $u\in X$. Then we claim that $X^{\prime}=X\setminus\{u\}$ is a general position set in $G-e$. If $v\notin X$, then we see by the same argument as above that $X^{\prime}$ is a general position set of $G-e$. It remains to consider the subcase when $v\in X$. Since ${\rm diam}(G)=2$, the only way that $X^{\prime}$ is not a general position set in $G-e$ would be when there is a shortest path $Q$ in $G-e$ containing $v$ and two other vertices of $X^{\prime}$. Since the path $Q$ does not contain the vertex $u$, $Q$ is also a shortest path in $G$, impossible. We can conclude that $X^{\prime}$ is a general position set of $G-e$ and therefore ${\rm gp}(G-e)\geq{\rm gp}(G)-1$.
Let $Y$ be an arbitrary ${\rm gp}$-set of $G-e$. Assume that $u,v\notin Y$. We claim that $Y$ is a general position set of $G$. Consider any two vertices $x,y$ from $Y$ and let $P$ be an arbitrary shortest $x,y$-path in $G$. Then the path $P$ does not contain the edge $uv$ in $G$ as ${\rm diam}(G)=2$, hence ${\cal P}_{G}(x,y)={\cal P}_{G-e}(x,y)$. We thus have that $Y$ is a general position set of $G$, and ${\rm gp}(G)\geq{\rm gp}(G-e)$. Assume now that $u\in Y$. Then we claim that $Y^{\prime}=Y\setminus\{u\}$ is a general position set of $G$. If $v\notin Y$, then similarly as above, $Y^{\prime}$ is a general position set of $G$. It remains to consider the case when $v\in Y$. Suppose on the contrary that there are three vertices from $Y^{\prime}$ such that they lie on a common shortest path $P^{\prime}$ in $G$. Then the path $P^{\prime}$ contains $v$ and two other vertices from $Y^{\prime}$. The path $P^{\prime}$ does not contain the edge $e$ and so $P^{\prime}$ is also a shortest path in $G-e$, a contradiction. Hence $Y^{\prime}$ is a general position set of $G$ and we can conclude that ${\rm gp}(G)\geq{\rm gp}(G-e)-1$.
To show that the lower bound is sharp, consider again fan graphs $F_{n}$, where $n=3k-1$, $k\geq 3$, see Fig. 6.
@@
Figure 6: Fan graph $F_{n}$ with $n=3k-1$.
By Proposition 2.4 we known that ${\rm gp}(F_{3k-1})=\lceil\frac{2(3k-1+1)}{3}\rceil=2k$, see Fig. 6 where the black vertices form a ${\rm gp}$-set of $F_{n}$. Let $uv$ be the edge of $F_{n}$ as shown in the figure. Then it is straightforward to check that all the black vertices but $v$ form a ${\rm gp}$-set of $F_{n}-e$, hence ${\rm gp}(F_{n}-e)={\rm gp}(F_{n})-1=2k-1$.
@@
7 Concluding remarks
In this paper we have explored how much the general position number of a graph can be affected by removing a vertex or by removing an edge. There are several problems that seem interesting to investigate in the future.
First, we have proved several lower and upper bounds and demonstrated that they are sharp. A challenging task remaining is to characterize the equality cases. It would also be of interest to determine sharper bounds for higher diameters. Two additional fundamental local graph operations which could be worth investigation with respect to the general position number are edge contraction and edge subdivision. Finally, the same line of research as in this paper could be done for different versions of the general position problem that appear in the literature, such as monophonic position sets.
@@ Acknowledgements
We would like to thank one of the reviewers for her/his accurate reading and many helpful tips. Among other things, the advice led to a shorter proof of Theorem 3.1.
This work was supported by the Slovenian Research Agency ARIS (research core funding P1-0297 and projects N1-0285, N1-0355).